Aleutian Company produces two products: Rings and Dings. They are manufactured in two departments: Fabrication and Assembly. Data for the products and departments are listed below.


Product
Number of
Units
Direct Labor Hours
Per Unit
Machine Hours
Per Unit
  Rings 970 6 8
  Dings 2,770 3 7

All of the machine hours take place in the Fabrication Department, which has estimated total factory overhead of $81,700. All of the labor hours take place in the Assembly Department, which has estimated total factory overhead of $70,000.

Aleutian Company uses the multiple production department factory overhead rate method. The Fabrication Department uses machine hours as an allocation base, and the Assembly Department uses direct labor hours.

The total factory overhead allocated per unit of Dings is

a.$4.51

b.$53.78

c.$40.56

d.$35.92
Answer #1

Total machine hours = Rings + Dings = (970 * 8) + (2,770 * 7) = 7,760 + 19,390 = 27,150

Fabrication department cost per machine hour = $81,700 / 27,150 = $3.01

Total direct labor hours = Rings + Dings = (970 * 6) + (2,770 * 3) = 5,820 + 8,310 = 14,130

Assembly department cost per labor hour = $70,000 / 14,130 = $4.95

Total factory overhead allocated per unit of Ding = (3 * $4.95) + (7 * $3.01)

= $14.85 + $21.07

= $35.92

The answer is D.